It is known that the electric force of repulsion b
Answer #1

Electrostatic repulsion, Fe = kq^2r^2
Gravitational attraction, Fg = Gm^2/r^2
Ratio, Fe/Fg = kq^2/Gm^2 = 4.1575*10^42